# Fully quantum algorithm for mesoscale fluid simulations with application to partial differential equations

### v3

@@ Abstract
Fluid flow simulations marshal our most powerful computational resources. In many cases, even this is not enough. Quantum computers provide an opportunity to speed up traditional algorithms for flow simulations. We show that lattice-based mesoscale numerical methods can be executed as efficient quantum algorithms due to their statistical features. This approach revises a quantum algorithm for lattice gas automata to eliminate classical computations and measurements at every time step. For this, the algorithm approximates the qubit relative phases and subtracts them at the end of each time step, removing a need for repeated measurements and state preparation (encoding). Phases are evaluated using the iterative phase estimation algorithm and subtracted using single-qubit rotation phase gates. This method delays the need for measurement to the end of the computation, avoiding expensive state preparation and measurement (SPAM) steps that are inefficient on modern hardware. We also demonstrate how the checker-board deficiency that the D1Q2 scheme presents can be resolved using the D1Q3 scheme. The algorithm is validated by simulating two canonical PDEs: the diffusion and Burgers’ equations on different quantum simulators. We find good agreement between quantum simulations and classical solutions of the presented algorithm.
quantum algorithms, lattice Boltzmann methods, fluid dynamics, CFD
@@ I Introduction
Solving partial differential equations (PDEs) is necessary for modeling most scientific problems. However, it is computationally expensive, especially for large domains and high-dimensional configuration spaces. For example, in computational fluid dynamics, the cost of performing a direct numerical simulation of turbulence increases with Reynolds number, representing the ratio of inertial to viscous effects in the flow, as $\mathcal{O}(\mathrm{Re}^{3})$ [1]. Such high computational expense motivates a broad category of work on achieving algorithmic speedups for PDE solvers.
@@
Quantum algorithms for solving PDEs have also gained attention. Harrow et al. [10] demonstrated the first quantum algorithm, HHL, for solving linear systems of equations with exponential speedup. Such improvements can be marshaled in the PDE space for solving discretized systems of equations via, for example, the method of lines. Childs and Liu [11] proposed a quantum algorithm for linear ordinary differential equations (ODEs) based on spectral methods, providing a global approximation to the solution using linear combinations of basis functions. Childs et al. [12] improved the complexity of quantum algorithms for linear PDEs using adaptive-order finite difference and spectral methods. Other approaches to solving PDEs were proposed by Berry [13]; for example, reducing an ODE system via discretization and solving it via the appropriate quantum algorithms. Gaitan [14] used this approach to solve the 1D Navier–Stokes equations for the flow through a convergent-divergent nozzle. However, these methods require deep circuits and many quantum gates to achieve results comparable to what one can solve on even a modern laptop. This proves prohibitive for most of the algorithms mentioned above but paints a bright future for larger, fault-tolerant quantum devices that can better reject gate noise.
@@
Other promising numerical methods include mesoscale methods, which operate between molecular and continuum scales. Lattice methods are a common example and are well suited for quantum computation because they are intrinsically statistical, resolving only samples or probabilities of the fictitious particles they comprise. Moreover, they are based on simple mathematical calculations and are suitable for parallel computation because interactions between lattice nodes are linear, and non-linearity enters during a local collision step. The lattice Boltzmann method (LBM) is a common approach that solves the Boltzmann transport equation. Mezzacapo et al. [20] developed the first quantum simulator following a lattice kinetic formalism to solve fluid dynamics transport phenomena. Todorova and Steijl [21] solved the collisionless Boltzmann equation on a quantum computer, inspired by quantum algorithms solving the Dirac equation. Budinski [22] proposed a novel quantum algorithm that solves the advection–diffusion equation by the LBM and extended it to solve the 2D Navier–Stokes equations using stream function–vorticity formulation [23]. This technique has a logarithmic scaling of qubits relative to lattice grid size. Itani and Succi [24] explored the Carleman linearization of the collision term of the lattice Boltzmann equation. Using this linearization technique, Itani et al. [25] proposed a quantum lattice Boltzmann algorithm where collision and streaming are achieved by unitary evolution.
@@
The original algorithm was designed for a specialized Type-II quantum computer, proposed in the early 2000s [27]. Such computers consisted of several small quantum computers connected by classical communication channels, allowing qubit shifts during the algorithm’s streaming step [27]. However, type II architectures are outdated, and we present an algorithm more suitable for modern architectures here that avoids some of the operations assumed to be efficient on older hardware.
The algorithm also requires estimating the distribution functions needed to re-prepare the quantum states at each time step. Re-preparation entails measuring each qubit at the end of each time step, a process that substantially increases computational time on current superconducting quantum computers.
Herein, we present an approach that overcomes the limitations of previous lattice-based quantum algorithms shown by Yepez [27, 28]. This improves the efficiency and scalability of previous QLGA techniques. Reading and re-prepare a quantum state is also exponentially expensive and is not scalable. This work seeks to mitigate these expense and scalability issues. The algorithm uses a quantum implementation of the streaming step, approximates the qubit relative phases, and subtracts them at the end of each time step, avoiding repeated encoding operations and delaying the need for measurement to the end of the computation.
We note that we do not interrogate the practicality of the presented algorithmic strategies on actual quantum devices, but rather on simulators. This is because the algorithm presented is not appropriate for solving practical fluid dynamics problems on quantum devices, for which, at present, there is perhaps no known algorithm. Instead, simulators help achieve validation and facilitate analysis. Specifically, we implement our algorithm in Qiskit [29] and XACC for the diffusion and Burgers’ equations.
This manuscript continues as follows. Section II describes the classical lattice gas automata and lattice Boltzmann method. Section III introduces the quantum counterpart of the classical LGA. Section IV presents the QLGA variants we propose, including quantum streaming, the D1Q3 scheme, and a measurement-avoiding algorithm. We validate these variants by testing them to solve the 1D diffusion and Burgers’ equations. Last, section V compares different quantum simulation strategies for the proposed algorithm using Qiskit and XACC.
@@ II Classical lattice methods
Figure 1: Illustration of the two main LBM operations: collision and streaming.
The main drawback of LGA is noise, which can be reduced by coarse-graining over large domains or time intervals but remains costly in computing time and memory. Instead, the LGA can be directly modeled at the mesoscopic scale with the Boltzmann equation [32, 33]. This approach is called the lattice Boltzmann method (LBM), which eliminates the LGA noise by replacing Boolean variables with continuous distributions, i.e., the mean occupation numbers, also called density distribution functions. We take this approach herein, which is thus a QLBM-type technique. The LBM approach replaces the collision rule with a continuous function called the collision operator. The distribution functions $f_{\alpha}$ ($0\leq f_{\alpha}\leq 1$) evolve in time according to the kinetic equation
@@
where $\bm{r}$ is the position of a lattice site, $\bm{c}_{\alpha}$ is the lattice velocity, $\Delta t$ is the time step, $t$ is time, $\Omega_{\alpha}$ is the collision operator, and ${\alpha}$ refers to the velocity discretization index. During one time step, the computation evolves through two processes: collision, where particles meet at the same lattice site, and their distribution functions are evaluated according to the collision operator $\Omega_{\alpha}$, and streaming, where particles shift to the neighboring sites following their lattice directions. Collision causes the relaxation of all local distribution functions to an equilibrium where $\Omega_{\alpha}[f_{\alpha}=f_{\alpha}^{\mathrm{eq}}]=0$.
@@ III Quantum lattice methods
Encoding and collision are performed in the quantum space, but the measurement step and classical streaming computation mean the quantum state must be re-encoded. The encoding step in QLA [26] scales linearly in the number of qubits required for encoding with respect to the size of the simulation. There are QLBM methods [22, 23] that scale logarithmically with respect to the size of the grid, though these entail a very large number of quantum gates (in many cases more than $10^{5}$ for even a small, toy problem) a trade-off that may be unappealing.
Figure 2: QLGA with classical streaming [26]. $R_{y}$ is a rotation gate by the angle $\theta_{\alpha}=2\arccos{\sqrt{1-f_{\alpha}(x_{i},t)}}$ for the qubit of state |$\ket{q_{\alpha}(x_{i},t)}$. The controlled $\sqrt{x}$ gate is a native IBM gate, decomposed as a two-qubit controlled phase rotation $P(\pi/2)$ gate between two Hadamard gates on the target qubit [35].
@@
Following Dirac’s bra-ket notation, the ket vector |$\ket{\Psi(x_{i},t)}$ represents the on-site state, i.e., at the lattice site of coordinate $x_{i}$, of dimension $2^{b}$. In the D1Q2 scheme, each lattice site comprises two particles, one moving to the right and the other to the left. QLGA maps the two particles to qubits, denoted by |$\ket{q_{\alpha}(x_{i},t)}$ where $\alpha=1,2$. The $\alpha$-th qubit encodes an occupancy probability $f_{\alpha}(x_{i},t)$ at lattice site $x_{i}$ and time $t$.
In the first step of the QLGA algorithm, qubits encode the occupancy probabilities via this expression:
@@
The on-site ket vector representing the quantum state at each lattice site $x_{i}$ is
@@
This step is implemented via the $R_{y}(\theta_{\alpha})$ rotation gate by the angle $\theta_{\alpha}=2\arccos{\sqrt{1-f_{\alpha}(x_{i},t)}}$ for the qubit of state |$\ket{q_{\alpha}(x_{i},t)}$. The relative phase of the qubits is set to zero in this step. Then, QLGA applies the collision locally to each lattice site, a unitary operation that relaxes the distribution functions to equilibrium. The post-collision state is
@@
Next, all qubits are measured, and the post-collision occupancy probabilities $f^{\prime}_{\alpha}(x_{i},t)$ are computed from measurement outcomes. The non-unitary measurement operation destroys the quantum superposition and entanglement that the collision step may have caused. The post-collision occupancy probability is
@@
where $n_{\alpha}$ are the number operators for $\alpha=1,2$.
Then, streaming is performed by shifting the qubits to their neighboring positions according to their right or left lattice directions. The post-streaming occupancy probabilities can be written as
@@
which is in the form of the classical lattice Boltzmann equation with the collision operator given by:
$\Omega_{\alpha}[f_{\alpha}(x_{i},t)]=\bra{\Psi(x_{i},t)}U^{\dagger}n_{\alpha}U-n_{\alpha}\ket{\Psi(x_{i},t)}$ (15)
Equilibrium distributions can be found by satisfying the equilibrium condition: $\Omega_{\alpha}[f_{\alpha}=f_{\alpha}^{\mathrm{eq}}]=0$. For the diffusion equation, we get equal equilibrium probabilities $f_{\alpha}^{\mathrm{eq}}=\rho/2$ for $\alpha=1,2$. For Burgers’ equation, the equilibrium distribution corresponds to $f_{\alpha}^{\mathrm{eq}}=\rho/2+e_{\alpha}/2(\sqrt{2}-\sqrt{1+(\rho-1)^{2}})$.
This streaming step was designed to run classically on Type-II quantum computers [27]. One time step of QLGA consists of the four operations described above. Unlike most quantum gate models, where measurement is performed at the end of the computation, QLGA requires measurement at each time step to extract the distribution functions, which are then used to encode the next time step. The recurrent decoherence at each lattice site is necessary to allow qubits to shift to neighboring sites classically. We next present a method for avoiding these unnecessary computations on modern quantum hardware.
@@ IV.1 Quantum streaming
The first revision to the QLGA algorithm performs the streaming step using quantum states, not classically. It can be implemented before the measurement step via the permutation function that applies a sequence of swap gates to shift the qubits to their neighbors. The corresponding quantum circuit is shown in fig. 3 for the diffusion and Burgers’ equations.
(a) QLBM circuit solving the diffusion equation
(b) QLBM circuit solving Burgers’ equation
@@
We validate the proposed variant by solving the 1D diffusion and Burgers’ equations using Gaussian and sinusoidal distributions as the initial condition. Periodic boundary conditions are used. We compare the two QLBM solutions, i.e., with classical streaming and quantum streaming, with the classical solution using the D1Q2 scheme and $26$ simulated qubits. The simulations were performed using Qiskit AerSimulator with the statevector simulation method, and the counts were sampled from the statevector using $10^{5}$ shots. The spatial domain consists of $N_{s}=13$ lattice sites corresponding to $26$ qubits. Results are shown after $10$ time steps.
@@
The results of fig. 4 demonstrate good agreement with the classical solution. This variant removes the need for classical communication between lattice sites and does not cause global entanglement, as swap gates merely interchange the qubit quantum states.
@@
In terms of complexity, we examine the relationship between the circuit depth and the system size by analyzing the different operations of the algorithm separately. The encoding step consists of applying one rotation gate simultaneously to each qubit. Thus, changing the grid resolution and the qubit number does not affect the circuit depth. The same unitary operator is applied to each pair of qubits for the collision step. Thus, the circuit depth scales as $\mathcal{O}(1)$ with the number of qubits $n$. For the streaming step implemented using the permutation function, the number of swap gates applied to one of the qubits increases linearly with the number of qubits. Therefore, the complexity of the algorithm is given by $\mathcal{O}(n)$. A D1Q3 scheme can also be implemented by using an extra qubit to create a smoother curve (fig. 9), as shown in the appendix A.
(a) Diffusion equation
(b) Burgers’ equation
@@
Figure 4: QLBM solutions to the diffusion equation (a) and Burgers’ equation (b) using the D1Q2 scheme. “Quantum - classical streaming” is associated with the classical streaming algorithm discussed in previous work [26],“Fully quantum” refers to the present algorithm, and “Classical” is the classical implementation of the lattice method.
IV.2 Measurement-avoiding QLBM

### v4

@@ Abstract
Abstract
Fluid flow simulations marshal our most powerful computational resources. In many cases, even this is not enough. Quantum computers provide an opportunity to speed up traditional algorithms for flow simulations. We show that lattice-based mesoscale numerical methods can be executed as efficient quantum algorithms due to their statistical features. This approach revises a quantum algorithm for lattice gas automata to reduce classical computations and state preparation at every time step. For this, the algorithm approximates the qubit relative phases and subtracts them at the end of each time step. Phases are evaluated using the iterative phase estimation algorithm and subtracted using single-qubit rotation phase gates. This method optimizes the quantum resource required and makes it more appropriate for near-term quantum hardware. We also demonstrate how the checkerboard deficiency that the D1Q2 scheme presents can be resolved using the D1Q3 scheme. The algorithm is validated by simulating two canonical PDEs: the diffusion and Burgers’ equations on different quantum simulators. We find good agreement between quantum simulations and classical solutions for the presented algorithm.
quantum algorithms, lattice Boltzmann methods, differential equations, fluid dynamics, CFD
@@ I Introduction
Solving partial differential equations (PDEs) is necessary for modeling most scientific problems. However, it is computationally expensive, especially for large domains and high-dimensional configuration spaces. For example, in computational fluid dynamics, the cost of performing a direct numerical simulation of turbulence increases with Reynolds number, representing the ratio of inertial to viscous effects in the flow, as $\mathcal{O}(\mathrm{Re}^{3})$ [1]. Such high computational expense motivates a broad category of work to achieve algorithmic speedups for PDE solvers.
@@
Quantum algorithms for solving PDEs have also gained attention. Harrow et al. [10] demonstrated the first quantum algorithm, HHL, for solving linear systems of equations with exponential speedup. Such improvements can be marshaled in the PDE space for solving discretized systems of equations via, for example, the method of lines. Childs and Liu [11] proposed a quantum algorithm for linear ordinary differential equations (ODEs) based on spectral methods, providing a global approximation to the solution using linear combinations of basis functions. Childs et al. [12] improved the complexity of quantum algorithms for linear PDEs using adaptive-order finite difference and spectral methods. Other approaches to solving PDEs were proposed by Berry [13]; for example, reducing an ODE system via discretization and solving it via the appropriate quantum algorithms. Gaitan [14] used this approach to solve the 1D Navier–Stokes equations for the flow through a convergent-divergent nozzle. However, these methods require deep circuits and many quantum gates to achieve results comparable to what one can solve on even a modern laptop. This limitation is prohibitive for most of the algorithms mentioned above but paints a bright future for larger, fault-tolerant quantum devices that can better reject gate noise.
@@
Other promising numerical methods include mesoscale strategies, which operate between molecular and continuum scales. Lattice methods are a common example and are well suited for quantum computation because they are intrinsically statistical, resolving only samples or probabilities of the fictitious particles they comprise. Moreover, they are based on simple mathematical calculations and are suitable for parallel computation because interactions between lattice nodes are linear, and non-linearity enters during a local collision step. The lattice Boltzmann method (LBM) is a common approach that solves the Boltzmann transport equation. Mezzacapo et al. [20] developed the first quantum simulator following a lattice kinetic formalism to solve fluid dynamics transport phenomena. Todorova and Steijl [21] solved the collisionless Boltzmann equation on a quantum computer, inspired by quantum algorithms solving the Dirac equation. Budinski [22] proposed a novel quantum algorithm that solves the advection–diffusion equation by the LBM and extended it to solve the 2D Navier–Stokes equations using stream function–vorticity formulation [23]. This technique has a logarithmic scaling of qubits relative to lattice grid size. Itani and Succi [24] explored the Carleman linearization of the collision term of the lattice Boltzmann equation. Using this linearization technique, Itani et al. [25] proposed a quantum lattice Boltzmann algorithm where collision and streaming are achieved by unitary evolution. Schalkers and Möller [26] proposes a space-time encoding method to implement unitary collision and streaming operators without linearization.
@@
The original algorithm was designed for a specialized Type-II quantum computer, proposed in the early 2000s [28]. Such computers consisted of several small quantum computers connected by classical communication channels, allowing qubit shifts during the algorithm’s streaming step [28]. Type II architectures no longer exist. Instead, we present an algorithm more suitable for current architectures that avoids some assumed-efficient operations on previous Type II devices. For example, many QLGA/LBM algorithms require estimating the distribution functions needed to re-prepare the quantum states at each time step. This process can be prohibitively expensive on current, and likely future, quantum devices.
Herein, we present an approach that overcomes the limitations of previous lattice-based quantum algorithms shown by Yepez [28, 29]. We focus on improving the efficiency and scalability of previous QLGA techniques. The cost of general quantum state preparation scales exponentially with respect to the qubit count [30]. This expensive step, assumed in previous quantum lattice-based methods using amplitude encoding, presents scaling issues for utility-scale simulations. Our work seeks to mitigate these expense and scalability issues. The proposed algorithm uses a quantum implementation of the streaming step, approximates the qubit relative phases, and subtracts them at the end of each time step, avoiding repeated encoding operations and delaying the need for measurement to the end of the computation.
We do not interrogate the practicality of the presented algorithmic strategies on actual quantum devices but rather on simulators. The algorithm presented herein is not appropriate for solving practical fluid dynamics problems, in full, on current quantum devices, for which there is perhaps no known algorithm at the time of writing. Instead, simulators help achieve validation and facilitate analysis. Specifically, we implement our algorithm in Qiskit [31] and XACC for the diffusion and Burgers’ equations.
This manuscript continues as follows. Section II describes the classical lattice gas automata and lattice Boltzmann method. Section III introduces the quantum counterpart of the classical LGA. Section IV presents the QLGA variants we propose, including quantum streaming, the D1Q3 scheme, and a restless QLBM algorithm. We validate these variants by testing them to solve the 1D diffusion and Burgers’ equations. Last, section V compares different quantum simulation strategies for the proposed algorithm using Qiskit and XACC.
@@ II Classical lattice methods
Figure 1: Illustration of collision and streaming steps for a single time-step of a D2Q5 scheme.
The main drawback of LGA is noise, which can be reduced via coarse-graining over large domains or time intervals but remains costly in computing time and memory. Instead, LGA can be directly modeled at the mesoscopic scale with the Boltzmann equation [34, 35]. This approach is called the lattice Boltzmann method (LBM), which eliminates the LGA noise by replacing Boolean variables with continuous distributions, i.e., the mean occupation numbers, also called density distribution functions. We take this approach herein, which is thus a QLBM-type technique. The LBM approach replaces the collision rule with a continuous function called the collision operator. The distribution functions $f_{\alpha}$ ($0\leq f_{\alpha}\leq 1$) evolve in time according to the kinetic equation
@@
where $\bm{r}$ is the position of a lattice site, $\bm{c}_{\alpha}$ is the lattice velocity, $\Delta t$ is the time step, $t$ is time, $\Omega_{\alpha}$ is the collision operator, and ${\alpha}$ refers to the velocity discretization index. During one time step, as illustrated in fig. 1, the computation evolves through two processes: collision, where particles meet at the same lattice site, and their distribution functions are evaluated according to the collision operator $\Omega_{\alpha}$, and streaming, where particles shift to the neighboring sites following their lattice directions. Collision causes the relaxation of all local distribution functions to an equilibrium where $\Omega_{\alpha}[f_{\alpha}=f_{\alpha}^{\mathrm{eq}}]=0$.
@@ III Quantum lattice methods
Encoding and collision are performed in the quantum space, but the measurement step and classical streaming computation mean the quantum state must be re-encoded. The encoding step in QLGA [27] scales linearly in the number of qubits required for encoding with respect to the size of the simulation. There are QLBM methods [22, 23] that scale logarithmically with respect to the size of the grid. However, these entail a prohibitively large number of quantum gates (in many cases more than $10^{5}$ for even a small, simplified problem) a trade-off that may be unappealing.
Figure 2: QLGA with classical streaming [27]. $R_{y}$ is a rotation gate by the angle $\theta_{\alpha}=2\arccos{\sqrt{1-f_{\alpha}(x_{i},t)}}$ for the qubit of state |$\ket{q_{\alpha}(x_{i},t)}$. The controlled $\sqrt{x}$ gate can be decomposed efficiently on the latest IBM Heron and Eagle processors, which consists of a two-qubit controlled phase rotation $P(\pi/2)$ gate between two Hadamard gates on the target qubit.
@@
Following Dirac’s bra-ket notation, the ket vector |$\ket{\Psi(x_{i},t)}$ represents the on-site state, i.e., at the lattice site of coordinate $x_{i}$, of dimension $2^{b}$. In the D1Q2 scheme, each lattice site comprises two particles, one moving to the right and the other to the left. The first step of the QLGA algorithm is mapping each particle to one qubit by encoding the occupancy probability $f_{\alpha}(x_{i},t)$ to $\alpha$-th qubit as
@@
This step is implemented via the $R_{y}(\theta_{\alpha})$ rotation gate with angle $\theta_{\alpha}=2\arccos{\sqrt{1-f_{\alpha}(x_{i},t)}}$ for the qubit of state |$\ket{q_{\alpha}(x_{i},t)}$. Then, the full quantum state at each lattice site $x_{i}$ is the tensor product state of two qubits
@@
The relative phase of the qubits is set to zero in this step. Then, QLGA applies the collision locally to each lattice site, a unitary operation that relaxes the distribution functions to equilibrium. The post-collision state is
@@
The next step is measuring all the $2N_{s}$ qubits associated with $N_{s}$ lattice sites. Notice the local nature of these measurements. The algorithm starts with a product state. The encoding (re-preparation) and collision steps are local operations inside each lattice site and, hence, do not change the structure of the product state. Thus, the number of measurements required in each time step scales linearly with the number of lattice sites. Then, the post-collision occupancy probabilities $f^{\prime}_{\alpha}(x_{i},t)$ are computed from measurement outcomes. The non-unitary measurement operation destroys the quantum superposition and entanglement that the collision step may have caused. The post-collision occupancy probability is
@@
where $n_{\alpha}$ are the number operators (observables that can be measured to count the number of particles) for $\alpha=1,2$.
Streaming is performed by shifting the qubits to their neighboring positions according to their right ($\alpha=1$) or left ($\alpha=2$) lattice directions. The post-streaming occupancy probabilities can be written as
@@
which is in the form of the classical lattice Boltzmann equation with the collision operator given by
$\Omega_{\alpha}[f_{\alpha}(x_{i},t)]=\bra{\Psi(x_{i},t)}(U^{\dagger}n_{\alpha}U-n_{\alpha})\ket{\Psi(x_{i},t)}.$ (11)
Equilibrium distributions can be found by satisfying the equilibrium condition: $\Omega_{\alpha}[f_{\alpha}=f_{\alpha}^{\mathrm{eq}}]=0$. For the diffusion equation, we get equal equilibrium probabilities $f_{\alpha}^{\mathrm{eq}}=\rho/2$ for $\alpha=1,2$ [37]. For Burgers’ equation, the equilibrium distribution corresponds to $f_{\alpha}^{\mathrm{eq}}=\rho/2+e_{\alpha}/2(\sqrt{2}-\sqrt{1+(\rho-1)^{2}})$ [38]. The equilibrium distributions for diffusion and Burgers’ equations differ due to the inherent differences in the underlying physical processes. The equilibrium state of the diffusion equation is characterized by maximum entropy and minimum energy. This equation results in a smooth, uniform distribution. Meanwhile, the equilibrium state of Burgers’ equation is characterized by the balance of convective and diffusive forces.
This streaming step was designed to run classically on Type-II quantum computers [28]. One time step of QLGA consists of the four operations described above. Unlike most quantum gate models, where measurement is performed at the end of the computation, QLGA requires measurement at each time step to extract the distribution functions, which are then used to encode the next step. Next, we present a method for improving these unnecessary computations that is more appropriate for modern quantum hardware.
@@ IV.1 Quantum streaming
The first revision to the QLGA algorithm performs the streaming step using quantum operations, not through classical post-processing. It can be implemented before the measurement step via the permutation function that applies a sequence of SWAP gates to shift the qubits to their neighbors. The corresponding quantum circuit is shown in fig. 3 for the diffusion and Burgers’ equations.
(a) (b)
Refer to caption Refer to caption
@@
We validate the proposed variant by solving the 1D diffusion and Burgers’ equations using Gaussian and sinusoidal distributions as the initial condition, respectively. Periodic boundary conditions are used. We compare the two QLBM solutions, i.e., with classical streaming and quantum streaming, with the classical solution using the D1Q2 scheme and $26$ simulated qubits. The simulations were performed using Qiskit AerSimulator with the statevector simulation method, and the counts were sampled from the statevector using $10^{5}$ shots. The spatial domain consists of $N_{s}=13$ lattice sites corresponding to $26$ qubits. Results are shown after $10$ time steps.
@@
The results of fig. 4 demonstrate good agreement with the classical solution. This variant removes the need for classical communication between lattice sites and does not cause global entanglement, as SWAP gates merely interchange the qubit quantum states. Since the quantum streaming does not change the local nature of product states, the total number of measurements required in each time step still scales linearly with the number of lattice sites.
@@
In terms of complexity, we examine the relationship between the circuit depth and the system size by analyzing the different operations of the algorithm separately. The encoding step consists of applying one rotation gate simultaneously to each qubit. Thus, changing the grid resolution and the qubit number does not affect the circuit depth. The same unitary operator is applied to each pair of qubits for the collision step. Thus, the circuit depth scales as $\mathcal{O}(1)$ with the number of qubits $n$. For the streaming step implemented using the permutation function, the number of SWAP gates applied to one of the qubits increases linearly with the number of qubits. Therefore, the complexity of the algorithm is given by $\mathcal{O}(n)$. A D1Q3 scheme can also be implemented by using an extra qubit to create a smoother curve (fig. 9), as shown in the appendix A. For a D1Q2 problem with a point source, particles do not remain in the original lattice site but propagate through the lattices over time. Therefore, a D1Q3 lattice scheme produces a smoother curve due to the non-zero distribution at one lattice site.
(a) (b)
Refer to caption Refer to caption
@@
Figure 4: QLBM solutions to (a) the diffusion equation and (b) Burgers’ equation using the D1Q2 scheme. “Quantum - classical streaming” corresponds to the classical streaming algorithm discussed in previous work [27], “Fully quantum” refers to the present algorithm using direct measurements, and “Classical” is the classical implementation of the lattice method. The small differences between “Quantum - classical streaming” and “Fully quantum” are due to finite sampling in quantum measurement.
IV.2 Restless QLBM